OpenStudy (anonymous):

Use implicit differentiation to find dy/dx if (x-2)^2+1/4y^2=1.

OpenStudy (bahrom7893):

2(x-2) + (1/2)y(dy/dx) = 0 (1/2)y(dy/dx) = -2(x-2) dy/dx = -4(x-2)/y

OpenStudy (anonymous):

thanks bahroom

OpenStudy (bahrom7893):

unless i made some dumb mistake.. and one o in my username lol

OpenStudy (anonymous):

my bad lol